OpenStudy (anonymous):

can you help me solve solve -x^3-3x^2+4=0 with steps

OpenStudy (tkhunny):

Have you considered the Rational Root Theorem?

OpenStudy (campbell_st):

well x = 1 looks like a solution to me you can use the factor theorem to check... and then a little division... so you can get the quadratic factor... then see if it can be factored

OpenStudy (anonymous):

how do i divide with an x^1 missing?

OpenStudy (campbell_st):

well if x = 1 is a solution then x - 1 is a factor... so if you don't know polynomial division or synthetic division... then use the rational root theorem

OpenStudy (anonymous):

i do know how but there is a remainder of 8. what do i do with that?

OpenStudy (tkhunny):

Put in a zero (0) as a place-holder.

OpenStudy (campbell_st):

factor out -1 then you have \[-1(x^3 + 3x^2 - 4) = 0\] it may be easier to see that x = 1 is a solution.. so (x -1) is a factor

hero (hero):

If x - 1 is a factor of -x^3 - 3x^2 + 4 then -x^3 + x^2 - 4x^2 + 4x - 4x+ 4 = 0 -x^2(x - 1) - 4x(x - 1) - 4(x - 1) = 0 (x - 1)(-x^2 - 4x - 4) = 0 -(x - 1)(x^2 + 4x + 4) = 0 -(x - 1)((x + 2)^2 = 0

hero (hero):

@chyu

hero (hero):

So now applying zero product property: x - 1 = 0 x + 2 = 0 And the values of x should be obvious at this point.